# Statistical Test on Diffusion Model-based Anomaly Detection by Selective Inference

### v3

@@ 4.3 Identification of $\mathcal{Z}$ by Parametric Programming
Over-conditioning causes a reduction in power due to excessive conditioning. A technique called Parametric Programming is utilized to explore all intervals along the one-dimensional line, resulting in (20). The truncated region $\mathcal{Z}$ can be represented using $\mathcal{Z}^{\mathrm{oc}}$ as
$\mathcal{Z}=\bigcup_{z\in\mathbb{R}\mid\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{X}_{\mathrm{obs}}}}\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z).$ (27)
The number of $\mathcal{Z}^{\mathrm{oc}}$ is obviously finite due to the finiteness of the number of polyhedra, but for practical purposes it grows exponentially, making it difficult to identify all of them. In many other SI studies, it is known that a search from $z_{\mathrm{min}}=(-10\sigma-|z_{\mathrm{obs}}|)$ to $z_{\mathrm{max}}(=10\sigma+|z_{\mathrm{obs}}|)$ is sufficient for practical use, where $\sigma$ is the standard deviation of the test statistic $T(\bm{X},\bm{X}^{\mathrm{ref}})$. An algorithm for calculating the selective $p$-value via Parametric Programming is summarized in Algorithm 2.
Algorithm 2 Selective $p$-value Computation by Parametric Programming
0:  $\bm{X}_{\mathrm{obs}},\bm{X}^{\mathrm{ref}}_{\mathrm{obs}},z_{\mathrm{min}},z_{\mathrm{max}}$ and $z_{\mathrm{obs}}:=T(\bm{X}_{\mathrm{obs}},\bm{X}^{\mathrm{ref}}_{\mathrm{obs}})$
1:  $\mathcal{Z}\leftarrow\emptyset$
2:  Obtain $\mathcal{M}_{\bm{X}_{\mathrm{obs}}}$ by (7)
3:  Compute $\bm{a}$, $\bm{b}$ by (20)
4:  $z\leftarrow z_{\mathrm{min}}$
5:  while $z<z_{\mathrm{max}}$ do
6:     Compute $\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z)$ and $\mathcal{M}_{\bm{X}(z)}$ by (26) for $z$
7:     if $\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{X}_{\mathrm{obs}}}$ then
8:        $\mathcal{Z}\leftarrow\mathcal{Z}\cup\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z)$
9:     end if
10:     $z\leftarrow\max{\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z)}+\gamma$, where $\gamma$ is small positive number.
11:  end while
12:  $p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|Z|>|z_{\mathrm{obs}}|\mid Z\in\mathcal{Z}\right)$
12:  $p_{\mathrm{selective}}$
@@ 5 Experiments
We compared our proposed methods (DMAD-test, DMAD-test-oc) with the other methods: naive method (naive), bonferroni correction (bonferroni), and permutation test (permutation) on type I error rate and power. The details of the methods for comparison are described in Appendix B. We executed the experiment on AMD EPYC 9474F processor, 48-core $3.6$GHz CPU and $768$GB memory.
@@ 5.1 Numerical Experiments
Experiments on the type I error rate and power were conducted with two types of covariance matrices: independent $\Sigma=I_{n}\in\mathbb{R}^{n\times n}$ and correlation $\Sigma=(0.5^{|i-j|})_{ij}\in\mathbb{R}^{n\times n}$. In the type I error rate experiments, we used only normal images. The synthetic dataset for normal images is generated to follow $\bm{X}=(X_{1},X_{2},\dots,X_{n})^{\top}\sim\mathcal{N}(\bm{0},\Sigma)$. We made 1000 normal images for $n\in\left\{64,256,1024,4096\right\}$. In the power experiments, we used only abnormal images. We generated 1000 abnormal images $\bm{X}=(X_{1},X_{2},\dots,X_{n})^{\top}\sim\mathcal{N}(\bm{\mu},\Sigma)$. The mean vector $\bm{\mu}$ is defined as $\mu_{i}=\Delta$ for all $i\in\mathcal{S}$, and $\mu_{i}=0$ for all $i\in[n]\backslash\mathcal{S}$, where $\mathcal{S}\subset[n]$ is the anomalous region with its position randomly chosen. The image size of the abnormal images was set to 4096, with signals $\Delta\in\left\{1,2,3,4\right\}$. In all experiments, we made the synthetic dataset for 1000 reference images to follow $\bm{X}^{\mathrm{ref}}=(X^{\mathrm{ref}}_{1},X^{\mathrm{ref}}_{2},\dots,X^{\mathrm{ref}}_{n})^{\top}\sim\mathcal{N}(\bm{0},\Sigma)$. The threshold was set to $\lambda=0.8$, and the kernel size of the averaging filter was set to $3$. All experiments were conducted under the significance level $\alpha=0.05$. The details of the architecture of the diffusion model we used are shown in Appendix C. The diffusion models were trained on the normal images from the synthetic dataset. The diffusion models were trained with $T=1000$ and the initial time step of the reverse process was set to $T^{\prime}=460$, and the reconstruction was conducted $5$ step samplings. The noise schedule $\beta_{1},\beta_{2},\dots,\beta_{T}$ was set to linear. In all experiment, we aim to generate new images through probabilistic sampling, $\eta$ was set to $1$.
In addition, we conducted robustness experiments against non-Gaussian noise. To evaluate the robustness of our proposed method’s performance, we used various non-Gaussian distribution families with different levels of deviation from the standard normal distribution $\mathcal{N}(0,1)$. We considered the following non-Gaussian distributions with a 1-Wasserstein distance $d\in\{0.01,0.02,0.03,0.04\}$ from $\mathcal{N}(0,1)$:
  • Skew normal distribution family (SND).
  • Exponentially modified gaussian distribution family (EMG).
  • Student’s $t$-distribution family ($t$-distribution).
Note that these distributions are standardized in the experiments. Figure 2 shows the probability density functions for distributions from each family, such that the $d$ is set to $0.04$. We run $1000$ trials for each distribution family and each 1-Wasserstein distance to calculate the type I error rate. The significance levels $\alpha$ were set to $0.05$ and $0.10$, and the image size was set to $256$.
Figure 2: Non-Gaussian distributions with $d=0.04$
Results.
@@
Figure 4 shows type I error rate comparison. The proposed methods DMAD-test and DMAD-test-oc can control the type I error rate at the significance level $\alpha$, and bonferroni can control the type I error rate below the significance level $\alpha$. In contrast, naive and permutation cannot control the type I error rate. Figure 4 shows power comparison. Since naive and permutation cannot control the type I error rate, their powers are not considered. Among the methods that can control the type I error rate, the proposed method has the highest power. DMAD-test-oc is over-conditioned and bonferroni is conservative because there are many hypotheses, so they have low power. Figure 5 shows the results of the robustness experiments. DMAD-test maintains good performance on the type I error rate for all the considered distribution families.

### v4

@@ 4.3 Identification of $\mathcal{Z}$
Over-conditioning causes a reduction in power due to excessive conditioning. A technique called parametric programming is utilized to explore all intervals along the one-dimensional line, resulting in (20). The truncation intervals $\mathcal{Z}$ can be represented using $\mathcal{Z}^{\mathrm{sub}}$ as
$\mathcal{Z}=\bigcup_{z\in\mathbb{R}\mid\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}}\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z).$ (23)
An algorithm for computing the selective $p$-value is summarized in Algorithm 2. Figure 2 illustrates the example of conditional sampling distribution. It shows how the subintervals determine the conditional sampling distribution.
Figure 2: Schematic illustration of the selective inference procedure for diffusion models. It shows how the image $\bm{X}(z)$ changes with $z$. The truncation intervals that yield the same anomalous region $\mathcal{M}_{\bm{X}(z)}$ as the observed anomalous region $\mathcal{M}_{\bm{x}}$ define the conditional sampling distribution. The $p_{\mathrm{selective}}$ denotes the proportion of probability mass within the truncation intervals.
Algorithm 2 Selective $p$-value by Parametric Programming
0:  $\bm{x},\bm{x}^{\mathrm{ref}}$
1:  Initialize $\mathcal{Z}\leftarrow\emptyset$
2:  Compute $\mathcal{M}_{\bm{x}}$, $\bm{a}$, $\bm{b}$, and $T(\bm{x},\bm{x}^{\mathrm{ref}})$ by (8), (21), (11)
3:  Initialize $z$ to a sufficiently small value
4:  while $z$ is not large enough do
5:     Compute $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$ and $\mathcal{M}_{\bm{X}(z)}$ by (34)
6:     if $\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}$ then
7:        $\mathcal{Z}\leftarrow\mathcal{Z}\cup\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$
8:     end if
9:     $z\leftarrow\max{\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)}+\gamma$, where $0<\gamma\ll 1$
10:  end while
11:  $p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X}(Z),\bm{X}^{\mathrm{ref}}(Z))|\geq|T(\bm{x},\bm{x}^{\mathrm{ref}})|\;\middle|\;Z\in\mathcal{Z}\right)$
11:  $p_{\mathrm{selective}}$
@@ 5 Experiments
We compared our proposed method (proposed) on type I error rate and power with the following methods (see Appendix C for details of those methods):
  • w/o-pp: An ablation study without the parametric programming technique described in §4.3. The $p$-value is computed using (22), with $\mathcal{Z}$ replaced by $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z^{\mathrm{obs}})$.
  • naive: This method is conventionally used in practice. The $p$-value is computed as (13).
  • bonferroni: This method uses the Bonferroni correction. Bonferroni correction is widely used for multiple testing correction.
  • permutation: This method uses the permutation test. A permutation test is widely used for non-parametric hypothesis testing.
The architecture of diffusion models used in all experiments is detailed in Appendix D. We executed all experiments on AMD EPYC 9474F processor, 48-core $3.6$GHz CPU and $768$GB memory.
@@ 5.1 Numerical experiments
Experiments on the type I error rate and power were conducted with two types of covariance matrices: independent $\Sigma=I_{n}\in\mathbb{R}^{n\times n}$ and correlation $\Sigma=(0.5^{|i-j|})_{ij}\in\mathbb{R}^{n\times n}$. In the type I error rate experiments, we used only normal images. The synthetic dataset for normal images is generated to follow $\bm{X}\sim\mathcal{N}(\bm{0},\Sigma)$. We made 1,000 normal images for $n\in\left\{64,256,1024,4096\right\}$. In the power experiments, we used only abnormal images. We generated 1,000 abnormal images $\bm{X}\sim\mathcal{N}(\bm{\mu},\Sigma)$. The mean vector $\bm{\mu}$ is defined as $\mu_{i}=\Delta$ for all $i\in\mathcal{S}$, and $\mu_{i}=0$ for all $i\in[n]\backslash\mathcal{S}$, where $\mathcal{S}\subset[n]$ is the anomalous region with its position randomly chosen. The image size of the abnormal images was set to 4096, with signals $\Delta\in\left\{1,2,3,4\right\}$. In all experiments, we made the synthetic dataset for 1,000 reference images to follow $\bm{X}^{\mathrm{ref}}\sim\mathcal{N}(\bm{0},\Sigma)$. The threshold was set to $\lambda=0.8$, and the kernel size of the averaging filter was set to $3$. All experiments were conducted under the significance level $\alpha=0.05$. The diffusion models were trained on the normal images from the synthetic dataset. The diffusion models were trained with $T=1000$ and the initial time step of the reverse process was set to $T^{\prime}=460$, and the reconstruction was conducted $5$ step samplings (see Appendix E for details). The noise schedule $\beta_{1},\beta_{2},\dots,\beta_{T}$ was set to linear. In all experiment, we generated normal-looking images through probabilistic sampling, $\eta$ was set to $1$. In addition, we conducted robustness experiments against non-Gaussian noise. The details of the robustness experiments are described in Appendix F.
Results
@@
Figure 4 shows the comparison results of type I error rates. The proposed methods proposed and w/o-pp can control the type I error rate at the significance level $\alpha$, and bonferroni can control the type I error rate below the significance level $\alpha$. In contrast, naive and permutation cannot control the type I error rate. Figure 4 shows the comparison results of powers. Since naive and permutation cannot control the type I error rate, their powers are not considered. Among the methods that can control the type I error rate, the proposed has the highest power. The ablation study w/o-pp is over-conditioned and bonferroni is conservative because there are many hypotheses, so they have low power.
@@
5.2 Real-world data experiments
Experimental setup
We conducted experiments using T2-FLAIR MRI brain scans from the Brain Tumor Segmentation (BraTS) 2023 dataset Karargyris et al. (2023); LaBella et al. (2023) and MVtec AD dataset Bergmann et al. (2019). The details of the experimental settings are described in Appendix G.
Results
Table 1: Type I error rate and power comparison on real-world datasets at the significance level of $\alpha(=0.05)$. The proposed and bonferroni methods control the type I error rate at or below $\alpha$ for each dataset, whereas naive fails to do so and is therefore unreliable. Each entry shows type I error rate (left) and power (right). Additionally, proposed achieves higher power than bonferroni. The Figures 9 and 10 show the results of the proposed and naive methods for the MRI images (BraTS) and MVTec AD dataset, respectively.
Dataset naive bonferroni proposed
Bottle 0.46/N.A 0.00/0.00 $\bm{0.04}$/$\bm{0.18}$
Cable 0.88/N.A 0.00/0.00 $\bm{0.02}$/$\bm{0.40}$
Grid 0.82/N.A 0.00/0.00 $\bm{0.06}$/$\bm{0.34}$
Transistor 0.86/N.A 0.00/0.00 $\bm{0.08}$/$\bm{0.28}$
BraTS (T2-FLAIR) 0.59/N.A 0.00/0.00 $\bm{0.05}$/$\bm{0.28}$
Table 1 shows the comparison of the type I error rate and power. The naive cannot control the type I error rate, while the proposed and bonferroni can control the type I error rate below the significance level $\alpha$. The proposed has higher power than bonferroni. In addition, the examples of the results for each image are shown in Appendix H.
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.527$
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$
(a) bottle (Left: Normal, Right: Anomaly)
$p_{\rm naive}:0.031$, $p_{\rm selective}:0.331$
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.003$
(b) Brain (Left: Normal, Right: Anomaly)
Figure 5: Real-world examples of the naive and proposed methods. For each sample, the $p_{\mathrm{selective}}$ is high (true negative) for normal images and low (true positive) for abnormal images. In contrast, the $p_{\mathrm{naive}}$ remains low for all images, indicating that it fails to control the type I error rate.
6 Conclusions, limitations and future works
We introduced the DAL-Test, a novel statistical procedure for anomaly localizations identified by a diffusion model. With the proposed DAL-Test, the false positive detection rate can be controlled with the significance level because statistical inference is conducted conditional on the fact that the anomalous regions are identified by using a diffusion model. We demonstrated that the DAL-Test has higher power than the bonferroni correction, the only existing method for controlling the false positive detection rate. However, the growing the size of the diffusion model also leads to increased computational demands. In future work, we will focus on improving the computational efficiency of the DAL-Test.
Acknowledgement
Appendix A Proofs
A.1 Proof of Theorem 1
Under the null hypothesis, the probability integral transform implies that
$p_{\mathrm{selective}}\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{X}}\right\}\sim\mathrm{Uniform}(0,1),$ (24)
and hence for any $\alpha\in(0,1)$,
$\mathbb{P}\left(p_{\mathrm{selective}}\leq\alpha\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{X}}\right\}\right)=\alpha,\;\forall\alpha\in(0,1).$ (25)
By marginalizing over the nuisance parameter $\mathcal{Q}_{\bm{x}}$, we have
$\mathbb{P}\left(p_{\mathrm{selective}}\leq\alpha\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{x}}\right)$ (26)
$=\int_{\mathbb{R}^{n}}\mathbb{P}_{\mathrm{H}_{0}}(p_{\mathrm{selective}}\leq\alpha\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{x}})\ \mathbb{P}_{\mathrm{H}_{0}}(\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{x}}\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}})d\mathcal{Q}_{\bm{x}}$
[…]
$=\alpha$
Therefore, we have
$\mathbb{P}_{\mathrm{H}_{0}}\left(p_{\mathrm{selective}}\leq\alpha\right)$ (27)
$=\sum_{\mathcal{M}_{\bm{x}}\in 2^{[n]}}\mathbb{P}_{\mathrm{H}_{0}}(\mathcal{M}_{\bm{x}})\ \mathbb{P}_{\mathrm{H}_{0}}(p_{\mathrm{selective}}\leq\alpha\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}})$
$=\alpha\sum_{\mathcal{M}_{\bm{x}}\in 2^{[n]}}\mathbb{P}_{\mathrm{H}_{0}}(\mathcal{M}_{\bm{x}})$
$=\alpha$
A.2 Proof of Theorem 2
The conditioning on $\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}$ implies
$\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\Leftrightarrow\left(I_{2n}-\frac{\tilde{\Sigma}\bm{\nu}\bm{\nu}^{\top}}{\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu}}\right)\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\Leftrightarrow\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}=\bm{a}+\bm{b}z,$ (28)
where $z=T(\bm{X},\bm{X}^{\mathrm{ref}})\in\mathbb{R}$. Hence,
$\left\{\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}\;\middle|\;\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right\}$ (29)
=$=\left\{\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}\;\middle|\;\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}=\bm{a}+\bm{b}z\right\}$
=$=\left\{\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}\;\middle|\;\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}\right\}$
=$=\left\{\bm{a}+\bm{b}z\;\middle|\;z\in\mathcal{Z}\right\},$
where $\bm{X}(z)=\bm{a}_{1:n}+\bm{b}_{1:n}z$. As a result
$T(\bm{X},\bm{X}^{\mathrm{ref}})\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right\}\sim\mathcal{TN}(0,\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu})$ (30)
Appendix B Calculating the subinterval $\mathcal{Z}^{\text{sub}}$ for diffusion models
We now show that a reconstruction error $\mathcal{E}$ via diffusion models can be expressed as a piecewise-linear function of $\bm{X}$. To show this, we see that both the forward process and reverse process of the diffusion model are piecewise-linear functions as long as we employ a class of U-Net described below. It is easy to see the piecewise-linearity of the forward process as long as we fix the random seed for $\epsilon_{t}$. To make the reverse process a piecewise-linear function, we employ a U-Net composed of piecewise-linear components such as ReLU activation functions and pooling layers. Then, $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$ is represented as a piecewise-linear function of $\mathbf{x}_{t}$. Moreover, since $f^{(t)}_{\theta}(\mathbf{x}_{t})$ in (4) is a composite function of $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$, it is also a piecewise-linear function. By combining them together, we see that $\mathbf{x}_{t-1}$ is written as a piecewise-linear function of $\mathbf{x}_{t}$. Therefore, the entire reconstruction process is a piecewise-linear function since it just repeats the above operation multiple times (see Algorithm 1). As a result, the entire mapping ${\mathcal{D}}(\mathbf{X})$ of the diffusion model is a piecewise-linear function of the input image $\bm{X}$. Moreover, since the averaging filter $\mathcal{F}$ and the absolute operation are also piecewise-linear functions, $|\mathcal{F}(\bm{X}-\mathcal{D}(\bm{X}))|(=\mathcal{E}(\bm{X}))$ is piecewise-linear. Exploiting this piecewise-linearity, the interval $\mathcal{Z}^{\text{oc}}$ can be computed. The following theorem tells that the subinterval $\mathcal{Z}^{\text{sub}}(\bm{a}+\bm{b}z)$ can be computed by solving a set of linear inequalities.
Theorem 3.
The piecewise-linear mapping $\mathcal{A}(\bm{X})$ can be expressed as a linear function of the input image $\bm{X}$ on each polyhedral region $\mathcal{P}_{k}$.
$\forall\bm{X}\in\mathcal{P}^{(k)},\enskip\mathcal{A}(\bm{X})=\bm{\delta}^{(k)}+\bm{\Delta}^{(k)}\bm{X},$ (31)
where $\bm{\delta}^{(k)}$ and $\bm{\Delta}^{(k)}$ for $k\in[K]$ are the constant vector and the coefficient matrix with appropriate dimensions for the $k$-th polyhedron, respectively. Using the notation in (22), since the input image $\bm{X}(z)$ is restricted on a one-dimensional line, each component of the output of $\mathcal{A}$ is written as
$\forall z\in[L^{(k^{\prime})}_{i},U^{(k^{\prime})}_{i}],\enskip\mathcal{A}_{i}(\bm{X}(z))=\kappa^{(k^{\prime})}_{i}+\rho^{(k^{\prime})}_{i}z,$ (32)
where $\kappa^{(k^{\prime})}_{i}\in\mathbb{R}$ and $\rho^{(k^{\prime})}_{i}\in\mathbb{R}$ for $k^{\prime}\in[K^{{}^{\prime}}_{i}]$ are the coefficient and the constant of the $k^{\prime}$-th interval $[L^{(k^{\prime})}_{i},U^{(k^{\prime})}_{i}]$, and $K^{{}^{\prime}}_{i}$ is the number of linear pieces of $\mathcal{A}_{i}$. For each $i\in[n]$, there exists $k^{\prime}\in[K^{{}^{\prime}}_{i}]$ such that $z\in[L^{(k^{\prime})}_{i},U^{(k^{\prime})}_{i}]$, then the inequality $\mathcal{A}_{i}(\bm{X}(z))\geq\lambda$, can be solved as
$(z)\begin{gathered}\coloneqq\begin{cases}\left[\max\left(L^{(k^{\prime})}_{i},(\lambda-\rho^{(k^{\prime})}_{i})/\kappa^{(k^{\prime})}_{i}\right),U^{(k^{\prime})}_{i}\right]&\text{if }\kappa^{(k^{\prime})}_{i}>0,\\ \left[L^{(k^{\prime})}_{i},\min\left(U^{(k^{\prime})}_{i},(\lambda-\rho^{(k^{\prime})}_{i})/\kappa^{(k^{\prime})}_{i}\right)\right]&\text{if }\kappa^{(k^{\prime})}_{i}<0.\end{cases}{}\end{gathered}$ (33)
By applying the above theorem, we denote the subinterval as
$\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)=\bigcap_{i\in[n]}\left[L_{i}(z),U_{i}(z)\right].$ (34)
Appendix C Comparison methods for numerical experiments
We compared our proposed method with the following methods:
  • proposed: The proposed method uses the parametric programming.
  • w/o-pp: The proposed method use over-conditioning (without parametric programming). The $p$-value is calculated as
    $p_{\mathrm{ablation}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X}(Z),\bm{X}^{\mathrm{ref}}(Z))|>|T(\bm{x},\bm{x}^{\mathrm{ref}})|\;\middle|\;Z\in\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z^{\mathrm{obs}})\right)$ (35)
  • naive: The naive method. This method uses a conventional $z$-test without any conditioning in (13).
    $p_{\mathrm{bonferroni}}=\min(1,2^{n}\cdot p_{\mathrm{naive}}).$
    • Calculate the observed test statistic $z^{\mathrm{obs}}$ by applying the observed image $\bm{x}$ to the diffusion model.
      $p_{\mathrm{permutation}}=\frac{1}{B}\sum_{b\in[B]}\bm{1}\{|z^{(b)}|>|z^{\mathrm{obs}}|\},$
      where $\bm{1}\{\cdot\}$ denotes the indicator function.
Appendix D Architecture of the U-Net
Figure 6 shows the architecture of the U-Net used in our experiments. The U-Net has three skip connections, and the Encoder and Decoder blocks. For image sizes $n\in\{64,256,1024,4096\}$, the corresponding spatial dimensions of images are $(1,d,d)$ where $d\in\{8,16,32,64\}$.
Figure 6: The architecture of the U-Net
Appendix E Accelerated reverse processes
Methods for accelerating the reverse process have been proposed in DDPM, DDIM Song et al. (2022). When taking a strictly increasing subsequence $\tau$ from $\{1,\cdots,T\}$, it is possible to skip the sampling trajectory from $\mathbf{x}_{\tau_{i}}$ to $\mathbf{x}_{\tau_{i-1}}$. In this case, equations (3) and (5) can be rewritten as
where
Appendix F Robustness of the proposed method
To evaluate the robustness of our proposed method’s performance, we used various non-Gaussian distribution families with different levels of deviation from the standard normal distribution $\mathcal{N}(0,1)$. We considered the following non-Gaussian distributions with a 1-Wasserstein distance $d\in\{0.01,0.02,0.03,0.04\}$ from $\mathcal{N}(0,1)$:
  • Skew normal distribution family (SND).
  • Exponentially modified gaussian distribution family (EMG).
  • Student’s $t$-distribution family ($t$-distribution).
Note that these distributions are standardized in the experiments. Figure 7 shows the probability density functions for distributions from each family, such that the $d$ is set to $0.04$. The significance levels $\alpha$ were set to $0.05$ and $0.10$, and the image size was set to $256$. Figure 8 shows the results of the robustness experiments.
DMAD-test maintains good performance on the type I error rate for all the considered distribution families.
Figure 7: Non-Gaussian distributions with $d=0.04$